The profit made by a business in 2018 was triple that made in 2019. The profit made in 2020 was R 3 million less than the profit made in 2019. If the total profit in these 3 years was R 67 million, what was the profit made in 2020? 


1
Expert's answer
2021-06-13T17:09:44-0400

Let the profit made in 2019 be R. X million.

Profit made in made in 2018 was R. 3X million.

Profit made in 2020 was R(x"-"3) million.


"3x+x+(x-3)=67"

"5x-3=67"

"5x=70"

"x=14"


Therefore, profit made in 2020="(14-3)=11"


=R.11 Million